OpenStudy (anonymous):

The equation, y = 10(1.0605)x, represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the value for 1911 (x = 0)? This is the prize money, in thousands. For instance, if you got 23, this represents $23,000. A. $10,605 B. $10,000 C. $0 Wouldnt the correct answer be A?

OpenStudy (anonymous):

I have to assume that you mean \[y = 10(1.0605)^x\]

OpenStudy (anonymous):

Yes sorry about that.

OpenStudy (anonymous):

If that is correct then let me ask you.. What is \(5^1\)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

What is \(5^0\)

OpenStudy (anonymous):

Ohhhhh then it would be nothing

OpenStudy (anonymous):

No

OpenStudy (anonymous):

What is \(5^2\)

OpenStudy (anonymous):

25

OpenStudy (anonymous):

and \(5^3\)

OpenStudy (anonymous):

125

OpenStudy (anonymous):

Ok, so lets look at it going backward: \[5^3 = 125\] \[5^2 = 25 = \frac{125}{5}\] \[5^1 = 5 = \frac{25}{5}\] \[5^0 =\ ? = \frac{5}{5}\]

OpenStudy (anonymous):

So what is \(5^0\)

OpenStudy (anonymous):

Then it would be... i think 1?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

nicely done!

OpenStudy (anonymous):

And this is the case for any number: \[b^0 = \frac{b}{b} = 1, \forall b \ne 0\]

OpenStudy (anonymous):

Ohhh.... Yea makes much more sense. I am so thankful for this cite!!!! My dad doesnt remember much from high school. :P So he often tells me things that are wrong... Thanks sooooo much! (:

OpenStudy (anonymous):

So if that's the case let's go back to your original problem

OpenStudy (anonymous):

What is \(10(1.0605)^0\)

OpenStudy (anonymous):

Alright. so then it would be 10,00

OpenStudy (anonymous):

Correct? (:

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

Nice work.

OpenStudy (anonymous):

Haha! Nope the credit goes to you (: Thanks a bunch!